OpenStudy (anonymous):

Find the equation, in standard form, of the line passing through the points (5,1) and (-2,3).

OpenStudy (anonymous):

with those points you could find the slope

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

y2-y1 over x2-x1

OpenStudy (anonymous):

oh ok and then use y=mx+b

OpenStudy (anonymous):

yes, you can use that equation to get standard form but you know how?

OpenStudy (anonymous):

i think so... the slope would be -2/7 then you put that into the m spot of the y=mx+b form right?

OpenStudy (anonymous):

yes and pick a point and plug it in the x and y to solve for b then change it into standard form

OpenStudy (anonymous):

ok cool so it would be y=-2/7-17/7?

OpenStudy (anonymous):

how would you write this in standard form?

OpenStudy (anonymous):

maybe one of us did something wrong. I got in slope intercept y = -1/5x + 17

OpenStudy (anonymous):

i keep getting 2/7 as the slope

OpenStudy (anonymous):

which points are you using?

OpenStudy (anonymous):

5,1 and -2,3

OpenStudy (anonymous):